# Ringdown signatures of Kerr black holes immersed in a magnetic field

### v2

@@ 1 Introduction
The characteristic exponential ringdown signature of the black hole remnant of a binary merger, as predicted by General Relativity, is readily apparent in the first LIGO-Virgo-KAGRA (LVK) event GW150914 [1]. The signature closely matches the expectations of the complex quasinormal mode (QNM) pseudospectrum following from a linearized analysis of Kerr black hole perturbations (for reviews, see [2, 3, 4]). While the fully dynamical post-merger behaviour is in many ways more complex than the naive setting for linear perturbation theory about a stationary black hole configuration, the data clearly indicates consistency to leading order with an underlying picture of dissipative ringdown, as anticipated due to the presence of a horizon. With the growing catalogue of binary merger events allowing more detailed use of the ringdown component to extract black hole parameters, the extent to which individual (linear) quasinormal modes can be isolated is coming under increasing scrutiny [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Presently, it is unclear if non-linearities arising from contamination of the merger itself could affect the experimental extraction of the waveform parameters. It is thought that these effects should not impact the extraction of the dominant $(2,2,0)$ QNM, but it is not clear how these effects could impact the reliable measurements of higher overtones [14, 21]. Care is needed in adding higher-order modes, given the potential to ‘overfit’ the data, which could lead to inaccurate values for the ringdown parameters.
@@
A distinct, but related, question is how the basic Kerr template for these modes could be impacted theoretically by the black hole environment. Early examples include scalar hair [22, 23], and black hole charge within the full Kerr-Newman geometry [24, 25]; see also [26, 27, 28]. While these generalizations were considered primarily as a way to test for any deviation from the Kerr spectrum, in this paper we will explore the impact of a more physical black hole environment, namely an asymptotically uniform magnetosphere as an approximation to the field structure sustained by an accretion disk. While an astrophysically relevant magnetosphere is not expected to produce large corrections to the Kerr QNM spectrum, and thus seems impractical to detect for black holes in the LVK window given the current level of precision [29], it may still provide a nontrivial ‘nuisance’ parameter that could impact the extraction of mass and spin within the Kerr waveform template. In particular, the QNM spectrum is known to be highly (and nonlinearly) sensitive to perturbations of the underlying potential, as characterized by the pseudospectrum for example [30, 31, 21]. Accordingly, our main goal will be to explore the impact of a magnetosphere on the modeling of ringdown from a merger via quasinormal modes.
@@
Black hole magnetospheres have been analyzed extensively, with perhaps the most notable motivation being their potential role as power sources for active galaxies, e.g. via the Blandford-Znajek (BZ) mechanism [32]. However, such energy extraction phenomena rely on the specifics of force-free electrodynamics in the black hole ergosphere, for which no analytic solutions are known beyond the perturbative regime of slow rotation. Thus, in considering ringdown, it is of interest to study cases where exact solutions do exist, and where the back-reaction on the black hole geometry can be fully taken into account. While the magnetized Kerr-Newman solution is technically the simplest, a more physical system requires an approximately uniform field away from the black hole. The regime where the energy density in the magnetic field is dominant is described by the Melvin solution [33], in which the spacetime asymptotics are also modified, which in general will completely alter the behaviour of the QNM spectrum. A Kerr black hole immersed in an asymptotically uniform magnetic field is described by the Ernst-Wild solution [34, 35], generalizing the static Ernst solution [36]. The QNM spectrum for this system was studied in [37] focusing on the regime of very large magnetic fields, where the Melvin-like asymptotics at large radius lead to the possibility of superradiance [38, 39, 40]. Our focus will instead be on the regime of small magnetic fields (relative to the black hole mass), which is more readily realized astrophysically.
@@
To make contact with astrophysical systems, where the magnetosphere is localized and the spacetime is asymptotically flat, we choose to expand perturbatively in the magnetic field strength $B$. On one hand, this is consistent with the small scale of observed astrophysical magnetic fields (in Planck units), but it also creates an important scale separation in the geometry. The non-asymptotically flat Melvin spacetime only becomes dominant outside the Melvin radius $r_{M}\sim 1/B\gg r_{+}$. Thus, for parametrically small $B$, the solution exhibits conventional Kerr-like asymptotics within a large intermediate interval in radius $r_{+}\ll r\ll r_{B}$. More technically, the equations for perturbations of the spacetime exhibit a transition from Kerr-like asymptotics at ${\cal O}(B^{2})$ to Melvin-like confining asymptotics at ${\cal O}(B^{4})$. As a result, at ${\cal O}(B^{2})$ the modes are primarily sensitive to the localized impact of the magnetic field near the Kerr horizon and photon sphere. Analyzing the spectrum of perturbations at this order also has the secondary benefit of technical simplicity. The analysis of linearized perturbations in the Ernst-Wild geometry is complicated by the fact that there is no known decoupling of the mode equations, which therefore need to be analyzed via complex numerical techniques [4, 41, 42, 43] even in the scalar case. However, working at ${\cal O}(B^{2})$ leads to simplifications, which we exploit in this exploratory study by also working perturbatively in the Kerr rotation parameter $a$. This allows for complete decoupling of the perturbation equations.
@@
We proceed to determine the scalar Ernst-Wild QNM spectrum as a function of the rotation parameter $a$ and the magnetic field $B^{2}$ for the low-lying modes most relevant for the ringdown signature. For simplicity, we choose to focus on scalar perturbations which, while unphysical, are known in the Kerr limit to behave similarly to the tensor modes. We quantify the impact of this assumption and the perturbative expansion in $B^{2}$ and $a$ in the next section, focusing on the application to mergers in the LVK catalogue. Using the QNM spectrum, we then construct an Ernst-Wild ringdown template and utilize the public LVK ringdown pipeline pyRing to assess its fit to a number of merger events in the catalogue. More precisely, we vary the magnetic field and explore its impact on extracting the primary parameters of mass and spin for the final black hole remnant. We conclude that while the effect is not large, it warrants consideration in a similar manner to the potential impact of nonlinear corrections to the ringdown.
@@ 2 Perturbations of the Ernst-Wild solution
Rotating black holes solutions immersed in asymptotically uniform magnetic fields are astrophysically well-motivated but necessarily complex, and it is remarkable that an analytic solution exists. The solution for a fully back-reacted Kerr black hole immersed in a magnetized Melvin geometry was initially obtained by Ernst and Wild [34], following the use of a Harrison transformation [44] on a seed Kerr-Newman metric with mass $M$, angular momentum $J$ (with rotation parameter $a\equiv J/M$), and charge $q$. The initial solution was subsequently corrected and analyzed in some detail [45, 46, 47, 48, 42].
One of the intriguing features of the Ernst-Wild solution is that for the full geometry to be electrically neutral, the black hole necessarily acquires an induced Kerr-Newman electric charge $q=-2aMB$ via the Wald mechanism [49]. Following [37], to second order in the spin parameter $a$, and taking $q=-2aMB$, the Ernst-Wild geometry takes the form [45, 34, 46, 47],
@@
where $\Lambda_{0}$ is used to remove the conical singularity at the poles [50], and
@@
where $B$ is the magnitude of the asymptotically uniform magnetic field strength. We will assume that the magnetic field surrounding the black hole is ‘weak’, namely that $BM\ll 1$, a condition that is readily satisfied by known astrophysical magnetospheres. As discussed above, this restriction also has the benefit of restoring Kerr-like asymptotics for perturbations when we formally truncate the geometry at ${\cal O}(\tilde{B}^{2})$, as illustrated in Appendix A. On restoring units, this limit takes the form [51, 52, 53],
@@ 2.1 Linearized perturbation analysis
Figure 1: The real (left) and imaginary (right) parts of the leading scalar $s=0$ and tensor $s=-2$ Kerr quasinormal mode frequencies $\omega_{s{\ell mn}}$ are shown as a function of the rotation parameter $\tilde{a}$ computed via the continued fraction technique.
@@
Following [38, 37], for geometries such as EW, the radial and angular parts of the equation can be separated at the expense of introducing couplings between the different multipoles $\ell$. We use the following ansatz
@@
Substitution of this ansatz into Eq. 8, and using the orthogonality of the spherical harmonics, allows the reduction to a system of nine coupled radial equations [37]. As noted above, we further assume that the magnetic field surrounding the black hole is weak, $\tilde{B}=BM\ll 1$, and truncating the system of coupled radial equations at ${\cal O}(\tilde{a}^{2},\tilde{B}^{2})$, we have
@@
    Perturbative analysis at ${\cal}(\tilde{a}\tilde{B}^{2})$: A perturbative expansion in $\tilde{B}$ is well-motivated physically on the grounds that astrophysical magnetospheres are not expected to approach $BM{\sim}1$. Moreover, the large $B$ asymptotics are those of the Melvin universe which exhibits normal (as opposed to quasinormal) modes, significantly impacting the spectrum analyzed in [37]. As described in Appendix A, the restriction to ${\cal O}(\tilde{B}^{2})$ restores the Kerr-like asymptotic boundary conditions, and ensures that the QNM only reflect the impact of the magnetic field near the black hole. A perturbative expansion in $\tilde{a}$ is less well-motivated physically, as the LVK catalogue centers on remnant black holes with $\tilde{a}\sim 0.7{-}0.8$. However, when combined with a perturbative expansion in $\tilde{B}$, this approximation allows for separable perturbations, and numerically even for $\tilde{a}\sim 0.7{-}0.8$, the QNM spectrum for the linearized geometry agrees quantitatively to $\sim$25with the exact Kerr modes. This comparison is illustrated in Fig. 2 (see also [54, 55]). The fractional errors are larger for the imaginary part than the real part of the modes. To formalize the expansion, we can introduce a small dimensionless parameter $\epsilon$, scaling $\tilde{a}\rightarrow\tilde{a}\epsilon^{2}$ and $\tilde{B}^{2}\rightarrow\tilde{B}^{2}\epsilon$, and then truncate the expansion at ${\cal O}(\epsilon^{3})$ which retains terms of ${\cal O}(\tilde{B}^{2},\tilde{a},\tilde{a}\tilde{B}^{2})$. Terms of ${\cal O}(\tilde{B}^{4})$ and ${\cal O}(\tilde{B}^{6})$, while parametrically allowed under this scaling regime, are small inside the Melvin radius for the $\tilde{B}$-range considered here and are excluded to ensure Kerr-like large radius asymptotics as described above.
@@
$\begin{split}ds^{2}&=\frac{1}{2r}(2M-r)(2+B^{2}r^{2}\sin^{2}\theta)dt^{2}\\ &+\frac{r}{4M-2r}(-2-B^{2}r^{2}+B^{2}r^{2}\cos^{2}\theta)dr^{2}\\ &+(r^{2}+\frac{1}{2}B^{2}r^{4}\sin^{2}\theta)d\theta^{2}\\ &+\frac{1}{4}r^{2}(4-B^{2}r^{2}+B^{2}r^{2}\cos 2\theta)\sin^{2}\theta\,d\phi^{2}\\ &-\frac{2a}{r}(2+B^{2}r^{2}(3+\cos^{2}\theta))\sin^{2}\theta\,dt\,d\phi.\end{split}$ (14)
The horizon radius $r_{+}$ and angular momentum $J$ reduce to the Kerr expressions at ${\cal O}(\tilde{a})$, while the horizon angular velocity is corrected at ${\cal O}(\tilde{a}\tilde{B}^{2}$) [37],
@@ 3 QNM spectrum via continued fractions
In this section the massless scalar quasinormal mode spectrum of the Ernst-Wild black hole will be investigated using Leaver’s well-known continued fraction technique [56, 57]. For this purpose, it is convenient to re-express (18) as an equation for $R_{\ell m}(r)=\bar{R}_{\ell m}(r)/r$ in terms of the original radial coordinate,
@@
×$\times(r-2M)^{-\sqrt{2M}\sqrt{\tilde{a}m\omega-2M\Omega^{2}-\rho}}$
×$\times e^{i\Omega r}\sum_{k=0}^{\infty}d_{n}\left(\frac{r-2M}{r}\right)^{k},$ (23)
@@
The recurrence relation given by Eq. 25 corresponds to a second-order difference equation, and there is one solution, referred to as the minimal solution [58], which leads to a convergent power series. For discrete values of $\omega=\omega_{s\ell mn}$ [^1] The spin subscript $s$ will be dropped when considering quasinormal modes arising from scalar perturbations. – the quasinormal frequencies – the series in Eq. 23 is absolutely convergent as $r\rightarrow\infty$ [56]. The ratio of successive $d_{n}$ is given by the following infinite continued fraction,
@@
Setting $n=0$ in Eq. 31 gives us back Eq. 30 provided that for $n<0,\ \alpha_{n}=\beta_{n}=\gamma_{n}=0$. Eq. 31 possesses an infinite number of roots and numerically the $n^{\mathrm{th}}$ inversion gives the $n^{\mathrm{th}}$ stable root [56], i.e. to determine the fundamental mode, one sets $n=0$ while to compute the $n^{\mathrm{th}}$ overtone, one sets $n$ equal to the desired overtone number.
@@
Thus, we implement Nollert’s technique for improved convergence [59], first solving asymptotically for the ‘remainder’ $R_{N}(\omega)$, namely the infinite continued fraction inferred from (31) for $n>N\gg 1$. $R_{N}$ satisfies the following recurrence relation for $N\gg 1$,

### v3

# Ringdown signatures in the Ernst-Wild geometry: modeling Kerr black holes immersed in a magnetic field

@@ 1 Introduction
The characteristic exponential ringdown signature of the black hole remnant of a binary merger, as predicted by General Relativity, is readily apparent in the first LIGO-Virgo-KAGRA (LVK) event GW150914 [LIGOScientific:2016lio]. The signature closely matches the expectations of the complex quasinormal mode (QNM) pseudospectrum following from a linearized analysis of Kerr black hole perturbations (for reviews, see [Kokkotas:1999bd, Berti:2009kk, Konoplya:2011qq]). While the fully dynamical post-merger behaviour is in many ways more complex than the naive setting for linear perturbation theory about a stationary black hole configuration, the data clearly indicates consistency to leading order with an underlying picture of dissipative ringdown, as anticipated due to the presence of a horizon. With the growing catalogue of binary merger events allowing more detailed use of the ringdown component to extract black hole parameters, the extent to which individual (linear) quasinormal modes can be isolated is coming under increasing scrutiny [CalderonBustillo:2020rmh, Capano:2021etf, Cotesta:2022pci, Isi:2022mhy, Finch:2022ynt, Cheung:2022rbm, Isi:2023nif, Wang:2023xsy, Carullo:2023gtf, Zhu:2023mzv, Baibhav:2023clw, Nee:2023osy, Correia:2023bfn, Qiu:2023lwo, Zhu:2024dyl, may2024nonlinear]. Presently, it is unclear if non-linearities arising from contamination of the merger itself could affect the experimental extraction of the waveform parameters. It is thought that these effects should not impact the extraction of the dominant $(2,2,0)$ QNM, but it is not clear how these effects could impact the reliable measurements of higher overtones [Zhu:2023mzv, cardoso2024physical]. Care is needed in adding higher-order modes, given the potential to ‘overfit’ the data, which could lead to inaccurate values for the ringdown parameters.
@@
A distinct, but related, question is how the basic Kerr template for these modes could be impacted theoretically by the black hole environment. Early examples include scalar hair [Bekenstein:1973mi, Bekenstein:1996pn], and black hole charge within the full Kerr-Newman geometry [PhysRevD.88.064048, Konoplya:2013rxa]; see also [Barausse:2014tra, cannizzaro2024impact, Rahman:2022fay]. While these generalizations were considered primarily as a way to test for any deviation from the Kerr spectrum, in this paper we will explore the impact of a more physical black hole environment, namely an asymptotically uniform magnetosphere as an approximation to the field structure sustained by an accretion disk. While an astrophysically relevant magnetosphere is not expected to produce large corrections to the Kerr QNM spectrum, and thus seems impractical to detect for black holes in the LVK window given the current level of precision [gupta2024possible], it may still provide a nontrivial ‘nuisance’ parameter that could impact the extraction of mass and spin within the Kerr waveform template. In particular, the QNM spectrum is known to be highly (and nonlinearly) sensitive to perturbations of the underlying potential, as characterized by the pseudospectrum for example [Nollert_1996, Jaramillo_2021, cardoso2024physical]. Accordingly, our main goal will be to explore the impact of a magnetosphere on the modeling of ringdown from a merger via quasinormal modes.
@@
Black hole magnetospheres have been analyzed extensively, with perhaps the most notable motivation being their potential role as power sources for active galaxies, e.g. via the Blandford-Znajek (BZ) mechanism [Blandford:1977ds]. However, such energy extraction phenomena rely on the specifics of force-free electrodynamics in the black hole ergosphere, for which no analytic solutions are known beyond the perturbative regime of slow rotation. Thus, in considering ringdown, it is of interest to study cases where exact solutions do exist, and where the back-reaction on the black hole geometry can be fully taken into account. While the magnetized Kerr-Newman solution is technically the simplest, a more physical system requires an approximately uniform field away from the black hole. The regime where the energy density in the magnetic field is dominant is described by the Melvin solution [Melvin:1965zza], in which the spacetime asymptotics are also modified, which in general will completely alter the behaviour of the QNM spectrum. A Kerr black hole immersed in an asymptotically uniform magnetic field is described by the Ernst-Wild solution [ErnstWild:1976Metric, Aliev:1989wz], generalizing the static Ernst solution [Ernst:1976mzr]. The QNM spectrum for this system was studied in [Brito:2014super] focusing on the regime of very large magnetic fields, where the Melvin-like asymptotics at large radius lead to the possibility of superradiance [Dolan_2013, Brito:2015oca, Brito_2020]. Our focus will instead be on the regime of small magnetic fields (relative to the black hole mass), which is more readily realized astrophysically.
@@
To make contact with astrophysical systems, where the magnetosphere is localized and the spacetime is asymptotically flat, we choose to expand perturbatively in the magnetic field strength $B$. On one hand, this is consistent with the small scale of observed astrophysical magnetic fields (in Planck units), but it also creates an important scale separation in the geometry. The non-asymptotically flat Melvin spacetime only becomes dominant outside the Melvin radius $r_{M}\sim 1/B\gg r_{+}$. Thus, for parametrically small $B$, the mode solutions for perturbations exhibit conventional Kerr-like asymptotics within a large intermediate interval in radius $r_{+}\ll r\ll r_{M}$. More technically, the equations for perturbations of the spacetime exhibit a transition from Kerr-like asymptotics at ${\cal O}(B^{2})$ to Melvin-like confining asymptotics at ${\cal O}(B^{4})$. As a result, at ${\cal O}(B^{2})$ the modes are primarily sensitive to the localized impact of the magnetic field near the Kerr horizon and photon sphere. Analyzing the spectrum of perturbations at this order also has the secondary benefit of technical simplicity. The analysis of linearized perturbations in the Ernst-Wild geometry is complicated by the fact that there is no known decoupling of the mode equations, which therefore need to be analyzed via complex numerical techniques [Konoplya:2011qq, Pani:2013pma, Brenner_2021, Ghosh:2023etd] even in the scalar case. However, working at ${\cal O}(B^{2})$ leads to simplifications, which we exploit in this exploratory study by also working perturbatively in the Kerr rotation parameter $a$. This allows for complete decoupling of the perturbation equations.
@@
We proceed to determine the scalar Ernst-Wild QNM spectrum as a function of the rotation parameter $a$ and the magnetic field $B^{2}$ for the low-lying modes most relevant for the ringdown signature. For simplicity, we choose to focus on scalar perturbations which, while unphysical, are known in the Kerr limit to behave similarly to the tensor modes. We quantify the impact of this assumption and the perturbative expansion in $B^{2}$ and $a$ in the next section, focusing on those modes of most interest in an application to mergers in the LVK catalogue. To gain some qualitative insight into the role the magnetic field strength $B$ in post-merger ringdown, we imagine excising the exterior region of the geometry beyond the Melvin radius and construct a scalar analogue of an Ernst-Wild ringdown template by extrapolating the QNM spectrum to larger values of $B$. Utilizing the public LVK ringdown pipeline pyRing, we assess its fit to a number of merger events in the catalogue. More precisely, we vary the magnetic field and explore its impact on extracting the primary parameters of mass and spin for the final black hole remnant. We conclude that while the effect is not large, it warrants consideration in a similar manner to the potential impact of nonlinear corrections to the ringdown.
@@ 2 Perturbations of the Ernst-Wild solution
Rotating black holes solutions immersed in asymptotically uniform magnetic fields are astrophysically well-motivated but necessarily complex, and it is remarkable that an analytic solution exists. The solution for a fully back-reacted Kerr black hole immersed in a magnetized Melvin geometry was initially obtained by Ernst and Wild [ErnstWild:1976Metric], following the use of a Harrison transformation [Harrison:1968HarrisonTransformations] on a seed Kerr-Newman metric with mass $M$, angular momentum $J$ (with rotation parameter $a\equiv J/M$), and charge $q$. The initial solution was subsequently corrected and analyzed in some detail [Gibbons:2013EW, Diaz:1985EW, Hedja:2015BHMagnetic, Astorino_2016, Brenner_2021].
One of the intriguing features of the Ernst-Wild solution is that for the full geometry to be electrically neutral, the black hole necessarily acquires an induced Kerr-Newman electric charge q=2aMBq=-2aMB via the Wald mechanism [Wald:1974np]. Following [Brito:2014super], to second order in the spin parameter $a$, and taking q=2aMBq=-2aMB, the Ernst-Wild geometry takes the form [Gibbons:2013EW, ErnstWild:1976Metric, Diaz:1985EW, Hedja:2015BHMagnetic],
@@
where $\Lambda_{0}$ is used to remove the conical singularity at the poles [Hiscock:1981EWFixed], and
@@
where $B$ is the magnitude of the asymptotically uniform magnetic field strength. We will assume that the magnetic field surrounding the black hole is ‘weak’, namely that $BM\ll 1$, a condition that is readily satisfied by known astrophysical magnetospheres. As discussed above, this restriction also has the benefit of restoring Kerr-like asymptotics for perturbations when we formally truncate the geometry at ${\cal O}(\tilde{B}^{2})$, as illustrated in Appendix A. On restoring units, this limit takes the form [piotrovich2010magnetic, Frolov:2010mi, Frolov_2012],
@@ 2.1 Linearized perturbation analysis
Figure 1: The real (left) and imaginary (right) parts of the leading scalar $s=0$ and tensor s=2s=-2 Kerr quasinormal mode frequencies $\omega_{s{\ell mn}}$ are shown as a function of the rotation parameter $\tilde{a}$ computed via the continued fraction technique.
@@
Following [Dolan_2013, Brito:2014super], for geometries such as EW, the radial and angular parts of the equation can be separated at the expense of introducing couplings between the different multipoles $\ell$. We use the following ansatz
@@
Substitution of this ansatz into Eq. 8, and using the orthogonality of the spherical harmonics, allows the reduction to a system of nine coupled radial equations [Brito:2014super]. As noted above, we further assume that the magnetic field surrounding the black hole is weak, $\tilde{B}=BM\ll 1$, and truncating the system of coupled radial equations at ${\cal O}(\tilde{a}^{2},\tilde{B}^{2})$, we have
@@
    Perturbative analysis at ${\cal}(\tilde{a}\tilde{B}^{2})$: A perturbative expansion in $\tilde{B}$ is well-motivated physically on the grounds that astrophysical magnetospheres are not expected to approach $BM{\sim}1$. Moreover, the large $B$ asymptotics are those of the Melvin universe which exhibits normal (as opposed to quasinormal) modes, significantly impacting the spectrum analyzed in [Brito:2014super]. As described in Appendix A, the restriction to ${\cal O}(\tilde{B}^{2})$ restores the Kerr-like asymptotic boundary conditions, and ensures that the QNM only reflect the impact of the magnetic field near the black hole. A perturbative expansion in $\tilde{a}$ is less well-motivated physically, as the LVK catalogue centers on remnant black holes with $\tilde{a}\sim 0.7{-}0.8$. However, when combined with a perturbative expansion in $\tilde{B}$, this approximation allows for separable perturbations, and numerically even for $\tilde{a}\sim 0.7{-}0.8$, the QNM spectrum for the linearized geometry agrees quantitatively to $\sim$25% with the exact Kerr modes. This comparison is illustrated in Fig. 2 (see also [Cardoso:2019mqo, Wagle:2021tam]). The fractional errors are larger for the imaginary part than the real part of the modes. To formalize the expansion, we can introduce a small dimensionless parameter $\epsilon$, scaling $\tilde{a}\rightarrow\tilde{a}\epsilon^{2}$ and $\tilde{B}^{2}\rightarrow\tilde{B}^{2}\epsilon$, and then truncate the expansion at ${\cal O}(\epsilon^{3})$ which retains terms of ${\cal O}(\tilde{B}^{2},\tilde{a},\tilde{a}\tilde{B}^{2})$. Terms of ${\cal O}(\tilde{B}^{4})$ and ${\cal O}(\tilde{B}^{6})$, while parametrically allowed under this scaling regime, are small inside the Melvin radius for the $\tilde{B}-$range considered here and are excluded to ensure Kerr-like large radius asymptotics as described above.
@@
$\begin{split}ds^{2}&=\frac{1}{2r}(2M-r)(2+B^{2}r^{2}\sin^{2}\theta)dt^{2}\\ &+\frac{r}{4M-2r}(-2-B^{2}r^{2}+B^{2}r^{2}\cos^{2}\theta)dr^{2}\\ &+(r^{2}+\frac{1}{2}B^{2}r^{4}\sin^{2}\theta)d\theta^{2}\\ &+\frac{1}{4}r^{2}(4-B^{2}r^{2}+B^{2}r^{2}\cos 2\theta)\sin^{2}\theta\,d\phi^{2}\\ &-\frac{2aM}{r}(2+B^{2}r^{2}(3+\cos^{2}\theta))\sin^{2}\theta\,dt\,d\phi.\end{split}$ (14)
[…]
@@ 3 QNM spectrum via continued fractions
In this section the massless scalar quasinormal mode spectrum of the Ernst-Wild black hole will be investigated using Leaver’s well-known continued fraction technique [Leaver:1985analyticQNMKerr, Leaver:PhysRevD.45.4713]. For this purpose, it is convenient to re-express (18) as an equation for $R_{\ell m}(r)=\bar{R}_{\ell m}(r)/r$ in terms of the original radial coordinate,
@@
$\times(r-2M)^{-\sqrt{2M}\sqrt{\tilde{a}m\omega-2M\Omega^{2}-\rho}}$
$\times e^{i\Omega r}\sum_{k=0}^{\infty}d_{n}\left(\frac{r-2M}{r}\right)^{k},$ (23)
@@
The recurrence relation given by Eq. 25 corresponds to a second-order difference equation, and there is one solution, referred to as the minimal solution [Gautschi1967computational], which leads to a convergent power series. For discrete values of $\omega=\omega_{s\ell mn}$ [^1] The spin subscript $s$ will be dropped when considering quasinormal modes arising from scalar perturbations. – the quasinormal frequencies – the series in Eq. 23 is absolutely convergent as $r\rightarrow\infty$ [Leaver:1985analyticQNMKerr]. The ratio of successive $d_{n}$ is given by the following infinite continued fraction,
@@
Setting $n=0$ in Eq. 31 gives us back Eq. 30 provided that for $n<0,\ \alpha_{n}=\beta_{n}=\gamma_{n}=0$. Eq. 31 possesses an infinite number of roots and numerically the $n^{\mathrm{th}}$ inversion gives the $n^{\mathrm{th}}$ stable root [Leaver:1985analyticQNMKerr], i.e. to determine the fundamental mode, one sets $n=0$ while to compute the $n^{\mathrm{th}}$ overtone, one sets $n$ equal to the desired overtone number.
@@
Thus, we implement Nollert’s technique for improved convergence [Nollert:1993qnm], first solving asymptotically for the ‘remainder’ $R_{N}(\omega)$, namely the infinite continued fraction inferred from (31) for $n>N\gg 1$. $R_{N}$ satisfies the following recurrence relation for $N\gg 1$,